Brianna's family spent $134 on 2 adult tickets and 3 youth tickets at an amusment park. Max's family spent $146 on 3 adult tickets and 2 youth tickets. What is the price of a youth ticket?
Found 2 solutions by ewatrrr, rfer:
Answer by ewatrrr(24785)


Hi,
This question Sets Up two equations with 2 unknowns
Let x and y represent the cost of the youth and adult ticket respectively
Question states***
3x + 2y = $134
2x + 3y = $146
using elimination to solve for x
9x +6y = 3*134 | Multiplying 1st equation by 3
-4x -6y = -2*146 | Multiplying 2nd equation by -2
5x = 3*134 - 2*146
5x = $110
x = $22, the cost of the youth ticket
CHECKING our Answer*** ($34 being the cost of the adult ticket) y = (134-3*22)/2
3*$22 + 2*$34 = $66 + $68 = $134

Answer by rfer(16322)

3y+2a=134 a=67-1.5y
2y+3a=146
--------------
2y+3(67-1.5y)=146
2y+201-4.5y=146
-2.5y=-55
y=22 tickets
------------------
2(22)+3a=146
3a=102
a=34 tickets